# Burgers equation from non-thermal stationary states in nearly-integrable gases

### v1

@@ Appendix B Molecular dynamics simulation algorithm
In the simulations shown in Fig. 5 we performed the following Monte Carlo algorithm with a fixed time step, which we denote $\Delta t$:
@@
Appendix C Additional formulas and derivations
In this Appendix we present additional computations showing: Markovianity of the GHD diffusion operator, existence and uniqueness of the solution to the Chapman-Enskog equation and verify that parameter $u$ transforms as a velocity while the diffusion coefficient is invariant under the Galilean transformations.
C.1 Diffusion operator and its markovianity
@@
$\mathfrak{D}_{\rho}$ $=\left(\left(\mathbf{1}-n\mathcal{T}\right)^{-1}\rho_{\mathrm{tot}}^{-1}\tilde{\mathfrak{D}}_{\rho}\rho_{\mathrm{tot}}^{-1}\left(1-n\mathcal{T}\right)\right),$ (121)
@@
$W(\lambda,\lambda^{\prime})=\rho(\lambda)f(\lambda)\left(T^{\mathrm{dr}}(\lambda,\lambda^{\prime})\right)^{2}\left|v_{\rho}(\lambda)-v_{\rho}(\lambda^{\prime})\right|,$ (123)
@@
$\frac{1}{2\pi}\int{\rm d}\lambda\mathfrak{D}_{\rho}(\lambda,\lambda^{\prime})=\int{\rm d}\lambda\left(\tilde{\mathfrak{D}}_{\rho}\rho_{\rm tot}^{-1}(1-n\mathcal{T})\right)(\lambda,\lambda^{\prime}),$ (125)
@@
C.2 Existence and uniqueness of the diffusion coefficient
@@
$\mathbf{1}\cdot\mathcal{A}^{\rm st}=(h_{1}^{\prime})^{\rm dr}v_{\rho}\cdot(1-n\mathcal{T})=h_{2}^{\prime}=h_{1}.$ (130)
In the second equality we used that $v_{\rho}=(h_{2}^{\prime})^{\rm dr}/(h_{1}^{\prime})^{\rm dr}$ and that $(1-n\mathcal{T})$ removes the dressing from the function that it acts upon. Recalling the definition (52) of $u$ we confirm the sought after relation. This proves that if the operator $\Gamma^{\rm st}$ is hermitian then there is always a solution to the equation specifying the diffusion coefficient. The operators $\Gamma^{\rm st}$ obtained from the Fermi’s golden rule or from RTA are hermitian which concludes the existence part of the proof.
@@
C.3 Galilean transformations
The Galilean transformations describe the effect of the boost, by a constant velocity $u$, on the system. We denote with bar quantities referring to the system moving with velocity $u$ with respect to the stationary one.
The quasi-particle distribution functions of the two systems obey $\bar{\rho}_{\rm p}(\lambda)=\rho_{\mathrm{p}}(\lambda-u)$. The parameter $u$ transforms as velocity under the boost,
$\bar{u}=u+u,$ (131)
as can be directly verified from its definition (52) and the transformation rule for the quasi-particles distributions. In a similar fashion we find transformation rules for the effective velocity and the flux Jacobian transform
$\bar{v}_{\rho}(\lambda)$ $=v_{\rho}(\lambda-u)+u,$ (132)
$\bar{\mathcal{A}}(\lambda,\lambda^{\prime})$ $=\mathcal{A}(\lambda-u,\lambda^{\prime}-u)+u.$
This implies that the solutions Chapman-Enskog equation (60) in the moving and in the stationary system are related by $\bar{a}(\lambda)=a(\lambda-u)$. The diffusion coefficient in the moving system is then
@@ C.3 Galilean transformations
Appendix D Linearization of the collision integral
@@
$V(k)=\mathcal{V}(k/k_{r}),\quad V(k)=V(-k).$ (134)
Constant $k_{r}$ is a characteristic momentum exchanged between the tubes, which we assume to be small. In the small momentum transfer limit, processes involving 1-particle hole excitation are dominant and higher processes contribute at higher powers of $k_{r}$ [52]. While it is possible to directly compute $\delta\mathcal{I}[\rho_{\mathrm{p}}]/\delta\rho_{\mathrm{p}}$, such path is not the easiest way to find the linear contribution to the collision integral. The task becomes much easier if instead of describing the perturbation with $\delta\rho_{\mathrm{p}}$ we switch the description to pseudo-energy perturbation. In the language of pseudo-energy perturbation we define the linearized collision integral as
$\tilde{\Gamma}^{\mathrm{st}}=\mathbb{F}_{1}^{\mathrm{st}}\cdot\frac{\delta\mathcal{I}_{0}}{\delta\epsilon}\Big|_{\rho_{\mathrm{p}}^{\mathrm{st}}}\cdot\left(\mathbf{1}-\mathcal{T}n\right)^{-1}.$ (135)
@@
$\langle f|\tilde{\Gamma}^{\mathrm{st}}g\rangle=\int{\rm d}\lambda f^{\mathrm{Dr,st}}\,\frac{\delta\mathcal{I}_{0}}{\delta\epsilon}\Big|_{\rho_{\mathrm{p}}^{\mathrm{st}}}\cdot g^{\mathrm{dr,st}}.$ (137)
Writing the matrix element we used the identity $\int{\rm d}\lambda f\cdot\mathbb{F}\cdot g=\int{\rm d}\lambda f^{\mathrm{Dr}}\cdot g$, which follows from the definition of the operator $\mathbb{F}$ [42]. Given a state in the non-evolving tube and a particle number in the evolving tube, stationary state in the evolving tube is fixed with the condition (34). As discussed in the main text, to linearize collision integral in the density imbalance limit we consider a perturbation $\rho_{\mathrm{p}}=\rho_{\mathrm{p}}^{\mathrm{st}}+\delta\rho_{\mathrm{p}}$, while keeping the state in the non-evolving tube ($\rho_{\mathrm{p}}^{(b)}$) fixed. Combining (29) with the dynamic structure factor (63) we write $\mathcal{I}_{0}$ in a form convenient for linearization
@@
$G$ $=V^{2}(k(p_{1},h_{1}))\rho_{\mathrm{p}}(h_{1})\,\rho_{\mathrm{h}}(p_{1})\,\rho_{\mathrm{p}}^{(b)}(p_{2})\rho_{\mathrm{h}}^{(b)}(h_{2})$
@@
$\times\delta(k(p_{1},h_{1})+k^{(b)}(p_{2},h_{2}))\,\delta(\omega(p_{1},h_{1})+\omega^{(b)}(p_{2},h_{2})),$
@@
Perturbing $\rho_{\mathrm{p}}$ we find $H+\delta H=\delta\epsilon(p_{1})-\delta\epsilon(h_{1})$, with no unity order contribution, since we linearize around stationary state. It follows that
@@
where we performed the trivial integral in $\lambda$. Next step is to change coordinates to the center of mass coordinates
@@
with the Jacobian of change of variables begin unity. Working in the small momentum limit, $\mu_{i}$ is small. In the CM variables Dressed momentum and energy are
@@
$V^{2}(k(p)-k(h))$ $=V^{2}\left(k^{\prime}(\lambda)\mu\right)+\mathcal{O}\left(\mu^{3}\right).$ (139)
@@
$G^{\mathrm{st}}$ $=\tilde{A}^{2}\left((k^{\mathrm{st}})^{\prime}(\lambda_{1})\mu_{1}\right)(k^{\mathrm{st}})^{\prime}(\lambda_{1})\,(k^{(b)})^{\prime}(\lambda_{2})$
@@
$\times\frac{1}{|\mu_{1}|}\delta\left(\mu_{2}+\mu_{1}\frac{(k^{\mathrm{st}})^{\prime}(\lambda_{2})}{(k^{(b)})^{\prime}(\lambda_{1})}\right)\delta\left(v^{\mathrm{st}}(\lambda_{1})-v^{(b)}(\lambda_{2})\right),$
where $v=\omega^{\prime}/k^{\prime}$ is the effective velocity (11). Changing variables under the Dirac delta we used that $k^{\prime}=2\pi\rho_{\mathrm{tot}}>0$. Plugging expanded $G^{\mathrm{st}}$ into (137) we find
$\langle f|$ $\tilde{\Gamma}^{\mathrm{st}}g\rangle=(2\pi)^{2}\int{\rm d}\lambda_{1}\,{\rm d}\mu_{1}\,{\rm d}\lambda_{2}\,V^{2}\left((k^{\mathrm{st}})^{\prime}(\lambda_{1})\mu_{1}\right)$
@@
where
$\zeta^{(b)}(v)=\frac{\rho_{\mathrm{p}}^{(b)}\,\rho_{\mathrm{h}}^{(b)}\,(k^{(b)})^{\prime}}{|(v^{(b)})^{\prime}|}\Bigg|_{\left(v^{(b)}\right)^{-1}\left(v\right)}\,\quad|\mathcal{V}^{2}|=\int{\rm d}z\,|z|\,\mathcal{V}^{2}(z).$
@@
Appendix E Additional numerical results and verifications
@@
Straightforward way to solve this set of equations is to truncate the basis to a finite dimension. Inverting the matrix $\tilde{\Gamma}^{\mathrm{st}}_{ij}$ we encounter a complication. Because of the particle conservation law (16) we have $\tilde{\Gamma}^{\mathrm{st}}_{0,j}=0$. Thus the matrix is rank deficient. From computations performed in the section C.2 follows $B^{\mathrm{st}}_{0}-u\,C^{\mathrm{st}}_{0}=0$. To be able to uniquely determine $\alpha_{i}$, one needs to use the constitutive relation (61), which fixes $\alpha_{0}=0$, resolving the rank deficiency problem. Higher components of $\alpha$ are fixed by the matrix equation
@@
We applied the presented method to the system of weakly coupled Lieb–Liniger tubes. Data presented in figure 3 were computed using the presented method.
We now turn to analyzing the function $\partial_{\varrho}\epsilon_{0}$ for the system of two Lieb–Liniger tubes. Suppose the environmental tube is in a boosted thermal state. Pseudo energy is then $\epsilon^{(b)}(\lambda)=-\beta\,\mu^{(b)}+\kappa\,k^{(b)}(\lambda)+\beta\,\omega^{(b)}(\lambda)$, where $\beta=1/T$ and $\kappa$ is a potential related with the tube’s velocity. In [52] it is shown that for the environmental tube in a boosted thermal state, the pseudo energy of the stationary state in the coupled tube is $\epsilon^{\mathrm{st}}(\lambda)=-\beta\,\mu+\kappa\,k^{\mathrm{st}}(\lambda)+\beta\,\omega^{\mathrm{st}}(\lambda)$. Consequently, the bare pseudo energy reads $\epsilon_{0}^{\mathrm{st}}(\lambda)=\beta\,\mu+\kappa\,h_{1}(\lambda)+\beta\,h_{2}(\lambda)$. To compute the bare pseudo-energy we differentiated (6) and used the relation between dressing operations (10). We see that the only free parameter is the chemical potential $\mu$, fixing number of particles in the evolving tube. Therefore, applying the chain rule, $\partial_{\varrho}\epsilon_{0}^{\mathrm{st}}=\frac{{\rm d}\mu}{{\rm d}\varrho}\partial_{\mu}\epsilon_{0}^{\mathrm{st}}=-\frac{{\rm d}\mu}{{\rm d}\varrho}\beta$. The derivative $\frac{{\rm d}\mu}{{\rm d}\varrho}$ can be further simplified
@@
For an athermal stationary state in the reference tube and finite interaction strength, the factor $\partial_{\varrho}\epsilon_{0}^{\mathrm{st}}$ has much more complicated structure. To compute it we begin by writing $\epsilon^{\mathrm{st}}(\lambda)$ as $(\epsilon^{\mathrm{st}})^{\prime}(\lambda)+\nu$, where $\nu$ is a constant parameter. For a fixed value of $\nu$, using (34) we are able to find $\epsilon^{\mathrm{st}}(\lambda)$ by solving the non-linear integral equation. Having solved equations for sufficiently many values of $\nu$ we are able to find $\varrho(\nu)$ and compute $\partial_{\varrho}\epsilon_{0}^{\mathrm{st}}(\lambda)$.
Finally let us comment on the estimation of the relaxation time. As discussed in the main text it is customary to estimate relaxation time of $\Gamma^{\mathrm{st}}$ as a gap between zero and the smallest non-zero eigenvalue. To numerically find the smallest non-zero eigenvalue of the operator $\Gamma^{\mathrm{st}}$, algorithm similar to the one outlined above might be used. It is straightforward to verify that if $\gamma$ is an eigenvalue of $\Gamma^{\mathrm{st}}$, then it satisfies the equation
@@
Figure 6: First smallest non-zero eigenvalues of $\tilde{\Gamma}_{\rho_{\mathrm{p}}^{\mathrm{st}}}$ as a function of $\varrho$ for the system in the same setup as in Fig. 3.

### v2

@@ Appendix C Molecular dynamics simulation algorithm
In the simulations shown in Fig. 5 we performed the following variation of discrete step Monte Carlo algorithm [84] with a fixed time step, which we denote $\Delta t$:
@@
Appendix D Additional formulas and derivations
In this Appendix we present additional computations showing: Markovianity of the GHD diffusion operator, existence and uniqueness of the solution to the ChE equation and verify that parameter $u$ transforms as a velocity while the diffusion coefficient is invariant under the Galilean transformations.
D.1 Diffusion operator and its Markovianity
@@
$\mathfrak{D}_{\rho}$ $=\left(\mathbf{1}-n\mathcal{T}\right)^{-1}\rho_{\mathrm{t}}^{-1}\tilde{\mathfrak{D}}_{\rho}\rho_{\mathrm{t}}^{-1}\left(\mathbf{1}-n\mathcal{T}\right),$ (127)
@@
$W(\lambda,\lambda^{\prime})=\rho_{\mathrm{p}}(\lambda)f(\lambda)\left(\mathcal{T}^{\mathrm{dr}}(\lambda,\lambda^{\prime})\right)^{2}\left|v_{\rho}(\lambda)-v_{\rho}(\lambda^{\prime})\right|,$ (129)
@@
$\frac{1}{2\pi}\int{\rm d}\lambda\mathfrak{D}_{\rho}(\lambda,\lambda^{\prime})=\int{\rm d}\lambda\left(\tilde{\mathfrak{D}}_{\rho}\rho_{\rm t}^{-1}(\mathbf{1}-n\mathcal{T})\right)(\lambda,\lambda^{\prime}),$ (131)
@@
D.2 Existence and uniqueness of the diffusion coefficient
@@
$\mathbf{1}\cdot\mathcal{A}^{\rm st}=(h_{1}^{\prime})^{\rm dr}v_{\rho}\cdot(\mathbf{1}-n\mathcal{T})=h_{2}^{\prime}=h_{1}.$ (136)
In the second equality we used that $v_{\rho}=(h_{2}^{\prime})^{\rm dr}/(h_{1}^{\prime})^{\rm dr}$ and that $(\mathbf{1}-n\mathcal{T})$ removes the dressing from the function that it acts upon. Recalling the definition (53) of $u$ we confirm the sought after relation. This proves that if the operator $\Gamma^{\rm st}$ is hermitian then there is always a solution to the equation specifying the diffusion coefficient. The operators $\Gamma^{\rm st}$ obtained from the Fermi’s golden rule or from RTA are hermitian which concludes the existence part of the proof.
@@
D.3 Galilean transformations
The Galilean transformations describe the effect of the boost, by a constant velocity $w$, on the system. We denote with bar quantities referring to the system moving with velocity $w$ with respect to the stationary one.
The quasiparticle distribution functions of the two systems obey $\bar{\rho}_{\rm p}(\lambda)=\rho_{\mathrm{p}}(\lambda-w)$. The parameter $u$ transforms as velocity under the boost,
$\bar{u}=u+w,$ (137)
as can be directly verified from its definition (53) and the transformation rule for the quasiparticles distributions. In a similar fashion we find transformation rules for the effective velocity and the flux Jacobian transform
$\bar{v}_{\rho}(\lambda)$ $=v_{\rho}(\lambda-w)+w,$ (138)
$\bar{\mathcal{A}}(\lambda,\lambda^{\prime})$ $=\mathcal{A}(\lambda-w,\lambda^{\prime}-w)+w.$
This implies that the solutions ChE equation (61) in the moving and in the stationary system are related by $\bar{a}(\lambda)=a(\lambda-w)$. The diffusion coefficient in the moving system is then
@@ D.3 Galilean transformations
Appendix E Linearization of the collision integral
@@
$\tilde{V}(k)=\mathcal{V}(k/k_{r}),\quad\tilde{V}(k)=\tilde{V}(-k).$ (140)
Constant $k_{r}$ is a characteristic momentum exchanged between the tubes, which we assume to be small. It can also be regarded as an inverse of characteristic range of the potential. In the small momentum transfer limit, processes involving 1 particle-hole excitation are dominant and higher processes contribute at higher powers of $k_{r}$ [68]. While it is possible to directly compute $\delta\mathcal{I}[\rho_{\mathrm{p}}]/\delta\rho_{\mathrm{p}}$, such path is not the easiest way to find the linear contribution to the collision integral. The task becomes much easier if instead of describing the perturbation with $\delta\rho_{\mathrm{p}}$ we switch the description to pseudoenergy perturbation. In the language of pseudoenergy perturbation we define the linearized collision integral as
$\tilde{\Gamma}^{\mathrm{st}}=\mathbb{F}_{1}^{\mathrm{st}}\cdot\frac{\delta\mathcal{I}_{0}}{\delta\epsilon_{0}}\Big|_{\rho_{\mathrm{p}}^{\mathrm{st}}}.$ (141)
@@
$\langle f|\tilde{\Gamma}^{\mathrm{st}}g\rangle=\int{\rm d}\lambda f^{\mathrm{Dr,st}}\,\frac{\delta\mathcal{I}_{0}}{\delta\epsilon_{0}}\Big|_{\rho_{\mathrm{p}}^{\mathrm{st}}}\cdot g^{\mathrm{st}}.$ (143)
Writing the matrix element we used the identity $\int{\rm d}\lambda f\cdot\mathbb{F}\cdot g=\int{\rm d}\lambda f^{\mathrm{Dr}}\cdot g$, which follows from the definition of the operator $\mathbb{F}$ [61]. Given a state in the non-evolving tube and a particle number in the evolving tube, stationary state in the evolving tube is fixed with the condition (34). As discussed in the main text, to linearize collision integral in the density imbalance limit we consider a perturbation $\rho_{\mathrm{p}}=\rho_{\mathrm{p}}^{\mathrm{st}}+\delta\rho_{\mathrm{p}}$, while keeping the state in the non-evolving tube $\rho_{\mathrm{p}}^{(b)}$ fixed. Combining (29) with the dynamic structure factor (64) we write $\mathcal{I}_{0}$ in a form convenient for linearization
@@
$G=$ $\tilde{V}^{2}(k(p_{1},h_{1}))\rho_{\mathrm{p}}(h_{1})\,\rho_{\mathrm{h}}(p_{1})\,\rho_{\mathrm{p}}^{(b)}(p_{2})\rho_{\mathrm{h}}^{(b)}(h_{2})$
@@
$\times\delta(k(p_{1},h_{1})+k^{(b)}(p_{2},h_{2}))$
$\times\delta(\omega(p_{1},h_{1})+\omega^{(b)}(p_{2},h_{2})),$
@@
We perturb $\rho_{\mathrm{p}}$ around stationary state and linearize. It follows that
@@
where we performed the trivial integral in $\lambda$. Next step is to change coordinates to the center of mass (CM) coordinates
@@
with the Jacobian of change of variables being unity. Working in the small momentum limit, $\mu_{i}$ is small. In the CM variables Dressed momentum and energy are
@@
$\tilde{V}^{2}(k(p)-k(h))$ $=\tilde{V}^{2}\left(k^{\prime}(\lambda)\mu\right)+\mathcal{O}\left(\mu^{3}\right).$ (145)
@@
$G^{\mathrm{st}}$ $=\tilde{V}^{2}\left((k^{\mathrm{st}})^{\prime}(\lambda_{1})\mu_{1}\right)(k^{\mathrm{st}})^{\prime}(\lambda_{1})\,(k^{(b)})^{\prime}(\lambda_{2})$
@@
$\times\frac{1}{|\mu_{1}|}\delta\left(\mu_{2}+\mu_{1}\frac{(k^{\mathrm{st}})^{\prime}(\lambda_{2})}{(k^{(b)})^{\prime}(\lambda_{1})}\right)$
$\times\delta\left(v^{\mathrm{st}}(\lambda_{1})-v^{(b)}(\lambda_{2})\right),$
where $v=\omega^{\prime}/k^{\prime}$ is the effective velocity (11). Changing variables under the Dirac delta we used that $k^{\prime}=2\pi\rho_{\mathrm{t}}>0$. Plugging expanded $G^{\mathrm{st}}$ into (143) we find
$\langle f|$ $\tilde{\Gamma}^{\mathrm{st}}g\rangle=(2\pi)^{2}\int{\rm d}\lambda_{1}\,{\rm d}\mu_{1}\,{\rm d}\lambda_{2}\,\tilde{V}^{2}\left((k^{\mathrm{st}})^{\prime}(\lambda_{1})\mu_{1}\right)$
@@
where $|\mathcal{V}^{2}|=\int{\rm d}z\,|z|\,\mathcal{V}^{2}(z)$ and
$\zeta^{(b)}(v)=\frac{\rho_{\mathrm{p}}^{(b)}\,\rho_{\mathrm{h}}^{(b)}\,(k^{(b)})^{\prime}}{|(v^{(b)})^{\prime}|}\Bigg|_{\left(v^{(b)}\right)^{-1}\left(v\right)}.$ (146)
@@
Appendix F Additional numerical results and verifications
@@
Straightforward way to solve this set of equations is to truncate the basis to a finite dimension. Inverting the matrix $\tilde{\Gamma}^{\mathrm{st}}_{ij}$ we encounter a complication. Because of the particle conservation law (16) we have $\tilde{\Gamma}^{\mathrm{st}}_{0,j}=0$. Thus the matrix is rank deficient. From computations performed in the section D.2 follows $B^{\mathrm{st}}_{0}-u\,C^{\mathrm{st}}_{0}=0$. To be able to uniquely determine $\alpha_{i}$, one needs to use the constitutive relation (62), which fixes $\alpha_{0}=0$, resolving the rank deficiency problem. Higher components of $\alpha$ are fixed by the matrix equation
@@
We applied the presented method to the system of weakly coupled Lieb–Liniger tubes. Data presented in Fig. 3 were computed using the presented method.
We now turn to analyzing the function $\partial_{\varrho}\epsilon_{0}$ for the system of two Lieb–Liniger tubes. Suppose the environmental tube is in a boosted thermal state. Pseudo energy is then $\epsilon^{(b)}(\lambda)=-\beta\,\mu^{(b)}+\kappa\,k^{(b)}(\lambda)+\beta\,\omega^{(b)}(\lambda)$, where $\beta=1/T$ and $\kappa$ is a potential related with the tube’s velocity. In [68] it is shown that for the environmental tube in a boosted thermal state, the pseudoenergy of the stationary state in the coupled tube is $\epsilon^{\mathrm{st}}(\lambda)=-\beta\,\mu+\kappa\,k^{\mathrm{st}}(\lambda)+\beta\,\omega^{\mathrm{st}}(\lambda)$. Consequently, the bare pseudo energy reads $\epsilon_{0}^{\mathrm{st}}(\lambda)=\beta\,\mu+\kappa\,h_{1}(\lambda)+\beta\,h_{2}(\lambda)$. To compute the bare pseudoenergy we differentiated (6) and used the relation between dressing operations (10). We see that the only free parameter is the chemical potential $\mu$, fixing number of particles in the evolving tube. Therefore, applying the chain rule, $\partial_{\varrho}\epsilon_{0}^{\mathrm{st}}=\frac{{\rm d}\mu}{{\rm d}\varrho}\partial_{\mu}\epsilon_{0}^{\mathrm{st}}=-\frac{{\rm d}\mu}{{\rm d}\varrho}\beta$. The derivative $\frac{{\rm d}\mu}{{\rm d}\varrho}$ can be further simplified
@@
For an athermal stationary state in the bath tube and finite interaction strength, the factor $\partial_{\varrho}\epsilon_{0}^{\mathrm{st}}$ has much more complicated structure. To compute it we begin by writing $\epsilon^{\mathrm{st}}(\lambda)$ as $(\epsilon^{\mathrm{st}})^{\prime}(\lambda)+\nu$, where $\nu$ is a constant parameter. For a fixed value of $\nu$, using (34) we are able to find $\epsilon^{\mathrm{st}}(\lambda)$ by solving the non-linear integral equation. Having solved equations for sufficiently many values of $\nu$ we are able to find $\varrho(\nu)$ and compute $\partial_{\varrho}\epsilon_{0}^{\mathrm{st}}(\lambda)$.
Finally, let us comment on the estimation of the relaxation time. As discussed in the main text it is customary to estimate relaxation time of $\Gamma^{\mathrm{st}}$ as a gap between zero and the smallest non-zero eigenvalue, see Fig.6. To numerically find the smallest non-zero eigenvalue of the operator $\Gamma^{\mathrm{st}}$, algorithm similar to the one outlined above might be used. It is straightforward to verify that if $\gamma$ is an eigenvalue of $\Gamma^{\mathrm{st}}$, then it satisfies the equation
@@
Figure 6: The smallest non-zero eigenvalues of $\tilde{\Gamma}^{\mathrm{st}}_{ij}$ as a function of $\varrho$ for the system in the same state as in panel (a) in Fig. 3.